Please help me solve this equation: sqrt (2) ( 1 - 2 * sqrt (10) )
Answer by stanbon(75887)

sqrt (2) ( 1 - 2 * sqrt (10) )
----
This is not an equation:
----
= sqrt(2) - 2*sqrt(20)
---
= sqrt(2) - 2*2*sqrt(5)
---
= sqrt(2) - 4sqrt(5)
========================
Cheers,
Stan H.
=======